# To Rely or Not to Rely? Evaluating Interventions for Appropriate Reliance on Large Language Models

### v2

@@ 3.4. Experimental Conditions
Control. To generate the baseline advice, we input the questions into GPT-4o through OpenAI’s API and record the responses directly. Participants receive a direct recommendation of the answer (such as “Option A” in Logical Reasoning and “120 chickpeas” in Numerical Analysis) in addition to a 2-4 sentence long explanation supporting the answer.
Reliance Disclaimer. We append the following static disclaimer statement, “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context”, while keeping the advice generated in Control the same. This disclaimer was generated through prompting the LLM to disclose its limitation as a language model to convey that its outputs should be used carefully. For consistency and to emulate real-world chatbot disclaimers like ChatGPT, we apply the same disclaimer to all advice.
@@
However, we acknowledge that the effectiveness of a static disclaimer may not persist in long-term deployment. Alarm fatigue describes the process of desensitization, where users learn to ignore unfounded warnings over time (Cvach, 2012). We hypothesize that disclaimers may be most effective for priming new users or in shorter interactions, like a tutoring chatbot that students only use occasionally. Some alternative ways to improve longevity of the effect are to dynamically insert disclaimers fluidly into the conversation—which makes them less repetitive and predictable (Bo et al., 2024), or to insert them only when error rates are particularly high (Cvach, 2012).
@@
Uncertainty Highlighting. When generating the Control advice, we also record the log probability of each token. To visually augment the baseline advice, we highlight the tokens with log probability under -1 in stronger red and indicate them to be uncertain, and the tokens under -0.5 in lighter pink and indicate them to be slightly uncertain. We selected these thresholds to balance highlighting uncertain text within the answer with minimizing visual overwhelm for the reader. We make manual adjustments to maintain coherence, such as extending highlighting to full words if they were only partially highlighted and removing highlighting on articles that do not impact semantic meaning. Participants were not told that the method of determining uncertainty is based on log probabilities. However, they were reminded of the definition of the highlighting colors in every survey page with the LLM advice. Our choice to highlight generation probability follows prior works like (Spatharioti et al., 2023), but we note that alternative methods of highlighting can offer better interpretation, such as edit probability (Vasconcelos et al., 2023a).
@@
Implicit Answer. This intervention semantically augments the content of the advice to avoid writing the answer directly, requiring more engagement to extract which answer the advice supports. For the Logical Reasoning questions, we instruct GPT-4o to never explicitly state the answer, and to use paraphrasing to generate an argument in support of the answer. For the Numerical Estimation questions, we manually removed the final calculation step from the advice generated in Control (as it was difficult to achieve this through prompting alone). This way, we ensure that participants see the intermediate steps of the calculation, and can arrive at the value advised by the LLM if they engage with the answer enough to calculate it themselves.
@@ 4. Results
We pre-registered our experimental design and analysis plan online 444https://aspredicted.org/8BD_WWN before conducting the experiments. Here, we report the results of the analyses described in Section 3, separately for LSAT and Numerical Estimation due to the inherent differences between these two tasks.
@@ 4.1. Participants
Performance outcomes for LSAT questions. For each intervention condition, there is a “transition matrix” showing the proportion of people who answered wrongly initially then answered wrongly with LLM advice, wrongly initially then answered correctly with LLM advice, correctly initially then answered wrongly with LLM advice, correctly initially then answered correctly with LLM advice.
Two figures labelled (a) and (b) depict percent error distributions and WoA distributions in Numerical Estimation. Subfigure (a) depicts the distribution of Stage 1 and Stage 2 percent error via four graphs: one for each intervention condition. Subfigure (b) depicts the distribution of weight of advice via four graphs: one for each intervention condition.
@@ 4.2. Task Performance
LSAT. Due to the high difficulty of the LSAT questions, participants were initially correct in only 36.0% of question instances, not far from the rate of random guessing (20%). Participants then either received correct or wrong LLM advice, equivalent to a 50% accuracy rate. Their second-stage (LLM-assisted) accuracy rate was 46.1%. Figure 4 shows transition matrices of the first- and second-stage performance outcomes in each condition. In the performance outcomes, there are no significant differences between Control and any of the interventions as evaluated by a test ( for all three interventions). Based on the collaborative performance, it is evident that while the LLM improved overall performance, the human-AI team does not perform optimally. In each question, if the human-LLM team relied on the team member that got the correct answer, then the Best Possible Performance would reach 62.1% accuracy (not 100%, because sometimes both the human and the LLM answered incorrectly). The gap between optimal and actual performance represents the performance lost to wrong reliance behaviour—either relying on wrong LLM advice or ignoring correct LLM advice.
@@
While unrelated to task performance, we further compute how much the participants relied on the advice for additional insights on their behaviour. The weight-of-advice (WoA) taken by each participant in each question is shown as distributions for each condition in Figure 5(b), where a WoA of 1 means fully adopting the LLM’s advice, a WoA of 0 means retaining the initial answer, and other values interpolate between these definitions. The distribution of WoA tends to be bimodal, with participants in Control, Reliance Disclaimer, and Uncertainty Highlighting predominantly favouring adoption of the LLM advice. Since Implicit Answer did not provide the final recommendation, and instead left it to the participants to apply the mathematical hints, the WoA distribution is more flattened, even with values in the negatives or exceeding 1. In our WoA analysis, we drop outlier values over a magnitude of 5 (such as switching from an initial estimate of 600 to a final estimate 4480 when the LLM advice is 544, which constitutes a WoA of -69.3). By the Mann-Whitney U test, Reliance Disclaimer’s WoA () and Implicit Answer’s WoA () are significantly smaller than Control’s WoA.
Three reliance figures labelled (a), (b), and (c). All of them have Relative LLM Reliance (RLR) on the x-axis, and Relative Self-Reliance (RSR) on the y-axis. Figure (a) depicts reference reliance outcomes. Figure (b) depicts reliance outcomes for LSAT questions. Figure (c) depicts reliance outcomes for Numerical Estimation questions.
@@ 4.3. Reliance Behaviour
Following the method of prior work (Schemmer et al., 2023), we drop the data instances where the participant’s initial answer agreed with the LLM advice for the reliance analysis, as it is not possible to measure deliberate reliance in those cases (see the third pathway in Figure 3). For Numerical Estimation, we do not use a strict definition of agreement, as it is unlikely for a participant’s initial estimate to exactly match the LLM advice. Instead, we adopt a heuristic of perceived agreement. For example, a participant may perceive a guess of 700 to be “pretty much the same” as an advice of 750. We use a conservative 10% boundary for perceived agreement, as participants were primed with the expectation that guesses within 10% of the true count would earn a bonus, making this range a reasonable proxy for perceived agreement.555We refine this heuristic with caveat that for smaller levels of advice, the boundary of perceived equivalence becomes very small. To address this, we include an additional 50 padding around the advice. As a robustness check, we also conduct reliance analysis using a 5% agreement boundary without additional padding and find similar significance trends to those reported. After excluding the instances of agreement, 87.1% of samples in the LSAT task and 88.8% in Numerical Estimation are retained.
@@
To answer RQ1, we compute the Relative LLM Reliance (RLR), Relative Self-Reliance (RSR), and Appropriate Reliance Ratio (ARR) across all data instances collected per task in each condition. For interpretation, RLR indicates the rate of reliance on the LLM when the advice is beneficial, and RSR indicates the rate of self-reliance when the LLM advice is harmful. ARR is the ratio of RLR to (RSR), where higher values signify more appropriate reliance on the LLM. In terms of reliance outcomes, over-reliance maps to low RSR (i.e. relying on the LLM when the advice is bad), under-reliance maps to low RLR (i.e. not relying on the LLM when the advice is good), and appropriate reliance maps to when both are high. Figure 6(a), reconstructed based on Figure 1 from (Schemmer et al., 2022), provides a visual reference for the relationship between reliance outcomes and RLR and RSR values. Table 1 displays all the metrics calculated for each condition, separately for the LSAT and Numerical Estimation tasks. For the interventions, there are also p-values indicating the significance of the metric difference between treatment and Control666This study followed a preregistered analysis plan using as the significance threshold. While no multiple-hypothesis corrections were applied (as per the preregistration), we acknowledge that adjusting for multiple tests could influence the significance level..
Control | Reliance Disclaimer | Uncertainty Highlighting | Implicit Answer | ||
LSAT | RLR | 0.66 | 0.71, p=.2 | 0.52, p=.1 | 0.58, p=.2 |
RSR | 0.44 | 0.56, p=.05 | 0.62, p=.01 (+) | 0.63, p=.002 (+) | |
ARR | 1.17 | 1.62, p=.04 (+) | 1.35, p=.2 | 1.54, p=.09 | |
Numerical Estimation (WoA >0.01) | RLR | 0.91 | 0.91, p=.4 | 0.93, p=.3 | 0.67, p<.001 (–) |
RSR | 0.15 | 0.22, p=.1 | 0.23, p=.1 | 0.43, p=.004 (+) | |
ARR | 1.07 | 1.17, p=.2 | 1.20, p=.1 | 1.18, r=.2 | |
Numerical Estimation (WoA >0.5) | RLR | 0.8 | 0.74, p=.2 | 0.8, p=.5 | 0.45, p=.17 |
RSR | 0.39 | 0.47, p=.1 | 0.42, p=.3 | 0.55, p=.007 (+) | |
ARR | 1.32 | 1.38, p=.4 | 1.37, p=.4 | 1.01, p=.1 | |
Numerical Estimation (WoA >0.99) | RLR | 0.47 | 0.39, p=.2 | 0.3, p=.02 (–) | 0.34, p=.05 |
RSR | 0.68 | 0.79, p=.02 (+) | 0.77, p=.05 | 0.67, p=.5 | |
@@
Depicts RLR, RSR, and ARR for each intervention category and each question type–showing Numerical estimation reliance metrics with WoA thresholds of 0.01, 0.5, and 0.99. The interventions with statistically significant reliance metric differences generally show increases in RSR or decreases in RLR. Only Reliance Disclaimer has a statistically significant increase in ARR.
LSAT. The visual representations of the metrics and their significances for LSAT are plotted in Figure 6(b), where Control is plotted in gray and we compare the interventions against it. We find that Reliance Disclaimer is the only intervention that has a significantly positive influence on the ARR, by reducing both over-reliance and under-reliance (). Uncertainty Highlighting and Implicit Answer both significantly improve self-reliance ( and , respectively), but at the cost of shifting overall reliance away from the LLM. Despite so, the reduction of RLR – signifying correct reliance on LLM – is not significant. Both interventions have higher ARR values than Control, but the difference is again not significant at our power level. These results indicate that a static, persistent disclaimer message may induce appropriate reliance more effectively than more specialized methods.
@@
Values and significance for RLR, RSS, and ARR are expressed at these three WoA threshold levels in both Table 1 and Figure 6(c). Unlike in the LSAT task, we do not observe any intervention to effectively improve appropriate reliance through increasing ARR. Across different WoA thresholds, Implicit Answer increased appropriate self-reliance the most (significant at WoA thresholds of 0.01 with and 0.5 at ). At the lowest WoA threshold, this also increases under-reliance through significantly reducing RLR ()—which is undesirable. This could be due to being forced to perform simple math calculations to see the advised number—which people may not do if they are satisfied with their own answer, or believe that extracting the advice requires more effort than the reward is worth. Uncertainty Highlighting also causes significant decrease in RLR at a high WoA threshold of 0.99 (), indicating that people are hesitant to fully use the advice. This corroborates Figure 5(b), which shows that Uncertainty Highlighting and Reliance Disclaimer both have reduced density of WoA values close to 1 compared to Control.
@@
LSAT confidence changes by reliance behaviour. Two double bar graphs labelled (a) and (b). Both depict average changes in confidence levels when answering LSAT questions. Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition. Average change in confidence is positive except for Correct Reliance Behaviour under the Highlighting condition. Only Highlighting has statistically different changes in confidence for right vs. wrong reliance behaviours, with positive confidence changes after wrong reliance behaviour and barely decreasing average confidence after right reliance behaviour. Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2. Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect). Both are grouped by condition. The average change in confidence is always positive. Only Disclaimer shows a statistically significant difference, with average change in confidence after a correct answer being significantly higher after appropriate reliance than after overreliance.
NE confidence changes by reliance behaviour, with a WoA threshold of 0.5. Two double bar graphs labelled (a) and (b). Both depict average changes in confidence levels when answering Numerical Estimation questions, using a WoA threshold of 0.5. Bar graph (a) shows average change in confidence after Wrong Reliance Behaviour, and average change in confidence after Correct Reliance Behaviour–both grouped by condition. Average change in confidence is always positive. Only Implicit Answer has statistically different changes in confidence for right vs. wrong reliance behaviours, with confidence changes close to 0 for wrong reliance behaviour and increases in confidence for right reliance behaviour. Bar graph (b) shows average change in confidence for questions where participants did not initially agree with LLM advice, but aligned with LLM advice in Stage 2. Changes in confidence are calculated separately for cases of appropriate reliance (when the LLM advice was correct), and cases of overreliance (where the LLM advice was incorrect). Both are grouped by condition. The average change in confidence is always positive. There are no statistically significant differences between confidence change after appropriate reliance and after overreliance.
@@ 4.4. Confidence Calibration and Perceptions
To address RQ2, we compare participants’ confidence in their pre-advice answer and their post-advice answer (Change in Confidence) at the question level, along with several post-experiment Likert scale ratings of their subjective perceptions. We compare cases where participants made correct reliance decisions with those where they made incorrect ones, and further focus on instances of positive versus negative LLM reliance (cases where participants actively relied on the LLM). Unless otherwise indicated, significance is calculated within each condition using the Mann-Whitney U test with as the significance threshold. Error bars in the bar plots of Figures 7-10 represent standard errors.
@@
LSAT. Figure 7(a) shows participants’ Change in Confidence in the LSAT task for correct and wrong reliance behaviours. The most salient takeaway is that on-aggregate, across all conditions (not explicitly shown in the plot), people get more confident when they over- and under-rely than when making the appropriate reliance decisions (). Within each condition, Uncertainty Highlighting is the only one where this observed effect is significant (), which indicates poor confidence calibration since it is undesirable for people to be more confident in wrong decisions.
@@
As an extension, Figure 7(b) specifically focuses on cases where people chose to rely on the LLM. Here, the confidence scores tend to increase more as people trusted the LLMs enough to use them. We observe that Reliance Disclaimer is the only condition that achieves calibrated confidence—the increase in confidence from using the LLM’s advice is lower when the participant over-relied than when they relied appropriately (). Though not significant, it is interesting to observe that the opposite trend can be demonstrated in some of the other interventions, where confidence increases more in cases of over-reliance (NL). This indicates that reliance interventions that operate by moderating users’ confidence in the LLM may inadvertently miscalibrate confidence with respect to appropriate reliance.
@@
Numerical Estimation. For this task, we repeat the confidence analyses and show them in Figure 8. For simplicity, reliance is determined based on the middle WoA threshold of 0.5, since the confidence outcomes between high and low WoA thresholds are qualitatively similar (the results for WoA thresholds of 0.01 and 0.99 are shown in Appendix D as Figures D.1-D.2). Implicit Answer is the only condition to achieve a significant difference in calibrating confidence changes to be higher in correct reliance outcomes (Figure 8(a)). However, no significant differences are seen in the behaviours when we select for instances that explicitly result in reliance on LLMs (Figure 8(b)), which may indicate that Implicit Answer is effective at reducing confidence for negative self-reliance—perhaps people recognize that they can benefit from the advice, but did not want to go through the effort of calculating the recommendation.
@@
Overall Post-Perception. Lastly, Figure 9 shows the mean Likert ratings per perceptions category by condition. The full subjective perception statements are shown in the figure’s labels (1=disagree, 7=agree). We calculate significance between the Control and the interventions using the Mann-Whitney U test with as threshold. Although all LLM conditions scored positively across the three categories, Control scored highest in subjective perceptions compared to all other the interventions. In particular, Uncertainty Highlighting is the only intervention that had significantly lower ratings in all categories compared to Control, indicating that our implementation may have induced too much uncertainty in the users. An alternative approach to moderate this effect may be to highlight highly certain tokens in green in addition to uncertain tokens in red. In addition, we find that the frequency of baseline LLM usage correlates positively with all three subjective perception measures – see Appendix E for the details of this analysis. The main implication here is that uncertainty calibration methods can have the undesirable outcome of inducing negative perceptions of accuracy and usefulness of the LLM, which increases overall algorithm aversion against LLM advice, rather than appropriate reliance.
A quadruple bar graph depicting post-experiment perceptions. Each colour represents a different intervention condition (Control, Disclaimer, Highlighting, Implicit Answer). All three post-experiment perceptions are depicted.
Two bar graphs labelled (a), and (b) showing average Time Ratios for each question type. The y-axes are both labelled Time Ratio. Both x-axes have Control, Disclaimer, Highlighting, and Implicit Answer as categories. Bar graph (a) shows the average Time Ratios for LSAT questions, for which Highlighting and Implicit Answer are close to 0.6 and Control and Disclaimer are close to 0.4. Bar graph (b) shows the average Time Ratios for Numerical Estimation questions, for which Implicit Answer is the highest at around 1.6, Disclaimer and Highlighting are close to 1.0, and Control is close to 0.8.

### v3

@@ 3.4. Experimental Conditions
Control. To generate the baseline advice, we input the questions into GPT-4o through OpenAI’s API and record the responses directly. Participants receive a direct recommendation of the answer (such as “Option A” in Logical Reasoning and “120 chickpeas” in Numerical Analysis) in addition to an explanation supporting the answer.
@@
Reliance Disclaimer. The following static disclaimer statement is appended to all advice generated in Control: “Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context”. This text was generated through prompting the LLM to disclose its limitation as a language model to convey that its outputs should be used carefully. For consistency and to emulate real-world chatbot disclaimers like ChatGPT, we apply the same disclaimer to all advice. However, we acknowledge that the effectiveness of a static disclaimer may not persist in long-term deployment. Alarm fatigue describes the process of desensitization, where users learn to ignore unfounded warnings over time (Cvach, 2012). We hypothesize that disclaimers may be most effective for priming new users or in shorter interactions, like a tutoring chatbot that students only use occasionally. Some alternative ways to improve longevity of the effect are to dynamically insert disclaimers fluidly into the conversation—which makes them less repetitive and predictable (Bo et al., 2024), or to insert them only when error rates are particularly high (Cvach, 2012).
@@
Uncertainty Highlighting. When generating the Control advice, we record the log probability of each token. Our choice to highlight generation probability follows prior works like (Spatharioti et al., 2023), but we note that alternative methods of highlighting can offer better interpretation, such as edit probability (Vasconcelos et al., 2023a). To visually augment the baseline advice, we highlight the tokens with log probability under -1 in stronger red and indicate them to be uncertain, and the tokens under -0.5 in lighter pink and indicate them to be slightly uncertain. We selected these thresholds to balance highlighting uncertain text within the answer with minimizing visual overwhelm for the reader. We make manual adjustments to maintain coherence, such as extending highlighting to full words if they were only partially highlighted and removing highlighting on articles that do not impact semantic meaning. Participants were not told that the method of determining uncertainty is based on log probabilities. However, they were reminded of the definition of the highlighting colors in every survey page with the LLM advice.
@@
Implicit Answer. This intervention semantically augments the content of the advice to imply the answer without writing it directly, requiring more engagement to extract the answer that the advice supports. For the Logical Reasoning questions, we append an instruction to the prompt to never explicitly state the answer’s text, and to use paraphrasing to generate an argument in support of the answer. For the Numerical Estimation questions, we manually removed the final calculation step from the advice generated in Control (as it was difficult to achieve this through prompting alone). In this way, we ensure that participants see the intermediate steps of the calculation, and can arrive at the value advised by the LLM if they engage with the answer enough to calculate it themselves.
@@ 4. Results
We pre-registered our experimental design and analysis plan online333https://aspredicted.org/8BD_WWN, then ran the experiment. Here, we report the analyses described in Section 3 separately for LSAT and Numerical Estimation due to the inherent differences between these two tasks.
@@ 4.2. Task Performance
LSAT. Due to the high difficulty of the LSAT questions, participants were initially correct in only 36.0% of question instances, not far from the rate of random guessing (20%). Participants then either received correct or wrong LLM advice, equivalent to a 50% accuracy rate. Their second-stage (LLM-assisted) accuracy rate was 46.1%. Figure 4 shows transition matrices of the first- and second-stage performance outcomes in each condition. In the performance outcomes, there are no significant differences between Control and any of the interventions as evaluated by a test ( for all three interventions). Based on the collaborative performance, it is evident that while the LLM improved overall performance, the human-AI team does not perform optimally. In each question, if the human-LLM team relied on the team member that got the correct answer, then the Best Possible Performance would reach 62.1% accuracy (not 100%, because sometimes both the human and the LLM answered incorrectly). The gap between optimal and actual performance represents the performance lost to inappropriate reliance behaviour—either relying on wrong LLM advice or ignoring correct LLM advice.
@@
While unrelated to task performance, we further compute how much the participants relied on the advice for additional insights on their behaviour. The weight-of-advice taken by each participant in each question is shown as distributions for each condition in Figure 5(b), where a WoA of 1 means fully adopting the LLM’s advice, a WoA of 0 means retaining the initial answer, and other values interpolate between these definitions. The distribution of WoA tends to be bimodal, with participants in Control, Reliance Disclaimer, and Uncertainty Highlighting predominantly favouring adoption of the LLM advice. Since Implicit Answer did not provide the final recommendation, and instead left it to the participants to apply the mathematical hints, the WoA distribution is more flattened, even with values in the negatives or exceeding 1. In our WoA analysis, we drop outlier values over a magnitude of 5 (such as switching from an initial estimate of 600 to a final estimate 4480 when the LLM advice is 544, which constitutes a WoA of -69.3). By the Mann-Whitney U test, Reliance Disclaimer’s WoA () and Implicit Answer’s WoA () are significantly smaller than Control’s WoA.
@@ 4.3. Reliance Behaviour
Following the methods of previous work (Schemmer et al., 2023), we drop the data instances where the participant’s initial answer agreed with the LLM advice for the reliance analysis, as it is not possible to measure deliberate reliance in those cases. For Numerical Estimation, we do not use a strict definition of agreement, as it is unlikely for a participant’s initial estimate to exactly match the LLM advice. Instead, we adopt a heuristic of perceived agreement. For example, a participant may perceive a guess of 700 to be “pretty much the same” as an advice of 750. We use a conservative 10% boundary for perceived agreement, as participants were primed with the expectation that guesses within 10% of the true count would earn a bonus, making this range a reasonable proxy for perceived agreement.444We refine this heuristic with caveat that for smaller levels of advice, the boundary of perceived equivalence becomes very small. To address this, we include an additional 50 padding around the advice. As a robustness check, we also conduct reliance analysis using a 5% agreement boundary without additional padding and find similar significance trends to those reported. After excluding the instances of agreement, 87.1% of samples in the LSAT task and 88.8% in Numerical Estimation are retained.
@@
To answer RQ1, we compute the Relative LLM Reliance (RLR), Relative Self-Reliance (RSR), and Appropriate Reliance Ratio (ARR) across all data instances collected per task in each condition. For interpretation, RLR indicates the rate of reliance on the LLM when the advice is beneficial, and RSR indicates the rate of self-reliance when the LLM advice is bad. ARR is the ratio of RLR to (RSR), where higher values signify more appropriate reliance on the LLM (less over-reliance and less under-reliance, or at least a positive trade-off such as much less over-reliance and only a small increase in under-reliance). In terms of reliance outcomes, over-reliance maps to low RSR (i.e. relying on the LLM when the advice is bad), under-reliance maps to low RLR (i.e. not relying on the LLM when the advice is good), and appropriate reliance maps to when both are high. Figure 6(a), reconstructed based on Figure 1 from (Schemmer et al., 2022), provides a visual reference for the relationship between reliance outcomes and RLR and RSR values. Table 1 displays all the metrics calculated for each condition, separately for the LSAT and Numerical Estimation tasks. For the interventions, there are also p-values indicating the significance of the metric difference between treatment and Control.
| Control | Reliance Disclaimer | Uncertainty Highlighting | Implicit Answer | |||||
| LSAT | RLR | 0.66 | 0.71, p=.2 | 0.52, p=.1 | 0.58, p=.2 | |||
| RSR | 0.44 | 0.56, p=.05 | 0.62, p=.01 (+) | 0.63, p=.002 (+) | ||||
| ARR | 1.17 | 1.62, p=.04 (+) | 1.35, p=.2 | 1.54, p=.09 | ||||
RLR | 0.91 | 0.91, p=.4 | 0.93, p=.3 | 0.67, p<.001 (–) | |||
| RSR | 0.15 | 0.22, p=.1 | 0.23, p=.1 | 0.43, p=.004 (+) | ||||
| ARR | 1.07 | 1.17, p=.2 | 1.20, p=.1 | 1.18, r=.2 | ||||
RLR | 0.8 | 0.74, p=.2 | 0.8, p=.5 | 0.45, p=.17 | |||
| RSR | 0.39 | 0.47, p=.1 | 0.42, p=.3 | 0.55, p=.007 (+) | ||||
| ARR | 1.32 | 1.38, p=.4 | 1.37, p=.4 | 1.01, p=.1 | ||||
RLR | 0.47 | 0.39, p=.2 | 0.3, p=.02 (–) | 0.34, p=.05 | |||
| RSR | 0.68 | 0.79, p=.02 (+) | 0.77, p=.05 | 0.67, p=.5 | ||||
@@
LSAT. The visual representations of the metrics and their significances for LSAT are plotted in Figure 6(b), where Control is plotted in gray and we compare the interventions against it. We find that Reliance Disclaimer is the only intervention that has a significantly positive influence on the ARR, by reducing both over-reliance and under-reliance (i.e. it is higher and further to the right than Control). Uncertainty Highlighting and Implicit Answer both significantly improve self-reliance, but at the cost of shifting reliance away from the LLM when following it would be appropriate (although this decrease is not statistically significant)—these interventions tend to reduce LLM reliance overall, both when the advice is bad and when it is good. Both interventions have higher ARR values than Control, but the difference is not significant at our power level. These results indicate that a static, persistent disclaimer message can induce appropriate reliance more effectively than more specialized methods.
@@
Values and significance for RLR, RSS, and ARR are expressed at these three WoA threshold levels in both Table 1 and Figure 6(c). Unlike in the LSAT task, we do not observe any intervention to effectively improve appropriate reliance. Across different WoA thresholds, Implicit Answer increased appropriate self-reliance the most (significant at WoA thresholds of 0.01 and 0.5). At the lowest WoA threshold, this also increases under-reliance—which is undesirable. This could be due to being forced to perform simple math calculations to see the advised number—which people may not do if they are satisfied with their own answer, or believe that extracting the advice requires more effort than the reward is worth. Uncertainty Highlighting also causes significant decrease in RLR at high WoA thresholds, indicating that people are hesitant to fully use the advice. This corroborates Figure 5(b), which shows that Uncertainty Highlighting and Reliance Disclaimer both have reduced density of WoA values close to 1 compared to Control.
@@ 4.4. Confidence Calibration and Perceptions
To address RQ2, we examine participant confidence in their pre-advice answer and their post-advice answer for each question, along with several post-experiment Likert scale ratings of their subjective perceptions. We compare the changes in confidence between cases where participants made correct reliance decisions with those where they made incorrect ones, and further focus on instances of positive versus negative LLM reliance (over-reliance). Unless otherwise indicated, significance is calculated within each condition using the Mann-Whitney U test with as the significance threshold. Error bars in the bar plots of Figures 7-10 represent standard errors.
@@
LSAT. Figure 7(a) shows participants’ LSAT question confidence changes for correct and wrong reliance behaviours. The most salient takeaway is that on-aggregate across all conditions (not explicitly shown in the plot), people get more confident when they over- and under-rely than when making the appropriate reliance decisions (). Within each condition, Uncertainty Highlighting is the only one where this observed effect is significant (), which indicates poor confidence calibration since it is undesirable for people to be more confident in wrong decisions.
@@
As an extension, Figure 7(b) specifically focuses on cases where people chose to rely on the LLM. Here, the confidence scores tend to increase more as people trusted the LLMs enough to use them. We observe that Reliance Disclaimer is the only condition that achieves calibrated confidence—the increase in confidence from using the LLM’s advice is significantly lower when the participant over-relied than when they relied appropriately (). Though not significant, it is interesting to observe that the opposite trend can be demonstrated in the other interventions, where confidence increases more in cases of over-reliance. This indicates that reliance interventions that operate by moderating users’ confidence in the LLM may inadvertently miscalibrate confidence with respect to appropriate reliance.
@@
Numerical Estimation. For this task, we repeat the confidence analyses and show them in Figure 8. For simplicity, reliance is determined based on the middle WoA threshold of 0.5, since the confidence outcomes between high and low WoA thresholds are qualitatively similar (the results for WoA thresholds of 0.01 and 0.99 are shown in Appendix D as Figures D.8-D.9). Implicit Answer is the only condition to achieve a significant difference in calibrating confidence changes to be higher in correct reliance outcomes. However, no significant differences are seen in the behaviours when we select for instances that explicitly result in reliance on LLMs, which may indicate that Implicit Answer is effective at reducing confidence for negative self-reliance—perhaps people recognize that they can benefit from the advice, but did not want to go through the effort of calculating the recommendation.
@@
Overall Post-Perception. Lastly, Figure 9 shows the mean Likert ratings per perceptions category by condition. The full subjective perception statements are shown in the figure’s labels (1=disagree, 7=agree). We calculate significance between the Control and the interventions using the Mann-Whitney U test with as threshold. Although all LLM conditions scored positively across the three categories, Control scored highest in subjective perceptions compared to all other the interventions. In particular, Uncertainty Highlighting is the only intervention that had significantly lower ratings in all categories compared to Control, indicating that our implementation may have induced too much uncertainty in the users. An alternative approach to moderate this effect may be to highlight highly certain tokens in green in addition to uncertain tokens in red. In addition, we find that the frequency of LLM usage correlates positively with all three subjective perception measures, but not with reliance behaviours – see Appendix E for details. The main implication here is that uncertainty calibration methods can have the undesirable outcome of inducing negative perceptions of accuracy and usefulness of the LLM, which increases overall algorithm aversion against LLM advice, rather than appropriate reliance.